Mary is three times as old as Jill. In five years, she will be twice as old as Jill. Find the present age of Mary.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Mary is three times as old as Jill. In five years, she will be twice as old as Jill. Find the present age of Mary.
----------------
M = 3J
M+5 = 2(J+5)
etc

Answer by ikleyn(52788)

.

Jill is x  years old

Mary is 3x years old.


In 5 years

    3x + 5 = 2*(x+5).


Simplify and find x

    3x + 5 = 2x + 10

    3x - 2x = 10 - 5

       x    =    5


ANSWER.  Mary is 3x = 3*5 = 15 years old now.

Solved.